### v1

# Bridge the Modality and Capacity Gaps in Vision-Language Model Selection

Abstract
Vision Language Models (VLMs) excel in zero-shot image classification by pairing images with textual category names. The expanding variety of Pre-Trained VLMs enhances the likelihood of identifying a suitable VLM for specific tasks. Thus, a promising zero-shot image classification strategy is selecting the most appropriate Pre-Trained VLM from the VLM Zoo, relying solely on the text data of the target dataset without access to the dataset’s images. In this paper, we analyze two inherent challenges in assessing the ability of a VLM in this Language-Only VLM selection: the “Modality Gap”—the disparity in VLM’s embeddings across two different modalities, making text a less reliable substitute for images; and the “Capability Gap”— the discrepancy between the VLM’s overall ranking and its ranking for target dataset, hindering direct prediction of a model’s dataset-specific performance from its general performance. We propose VLM Selection With gAp Bridging (Swab) to mitigate the negative impact of these two gaps. Swab first adopts optimal transport to capture the relevance between open-source datasets and target dataset with a transportation matrix. It then uses this matrix to transfer useful statistics of VLMs from open-source datasets to the target dataset for bridging those two gaps and enhancing the VLM’s capacity estimation for VLM selection. Experiments across various VLMs and image classification datasets validate Swab’s effectiveness.
Machine Learning, ICML
1 Introduction
Vision-Language Models (VLMs) (Radford et al., 2021; Jia et al., 2021; Singh et al., 2022; Yuan et al., 2021) have demonstrated impressive image-text matching ability. One notable application of VLMs is zero-shot image classification (Radford et al., 2021; Menon & Vondrick, 2023; Ge et al., 2023; Mao et al., 2023), where VLMs are leveraged to generate image classifiers using only class names directly. This zero-shot approach has shown considerable success in image classification, particularly in scenarios with scarce or no training images (Ma et al., 2021; He et al., 2023).
Figure 1: Paradigm of Language-Only VLM Selection (LOVM). The model selection algorithm uses two types of data, including the open-source datasets (which have image and text data) and the text data from the target dataset, to predict the VLM’s absolute or relative performance on a target dataset. It then selects the most appropriate VLM based on the predicted performance.
Despite the success of VLM in image classification, the performance of a VLM may vary substantially according to the datasets and domains (Fang et al., 2022), making it challenging to use a single model to handle all tasks. Fortunately, many open-source VLMs are available (Ilharco et al., 2021), and these VLMs form a vast VLM Zoo. With different architectures, pre-training datasets, or training methods, these VLMs have different strengths. The diverse pre-trained VLMs increase the likelihood of pinpointing at least one VLM that excels in a given target dataset in most cases. [^1] Throughout this paper, the term “VLM” specifically refers to a pre-trained VLM. So one solution for zero-shot image classification is identifying the most suitable VLMs in the zoo for a target dataset without access to the dataset’s images. This VLM selection is termed as “Language-Only VLM Selection” (LOVM) (Zohar et al., 2023), and the paradigm is illustrated in Figure 1.
Two key types of information are available for LOVM. One is the target dataset’s text data, i.e., names of the target classes. The other is the open-source datasets, collected in the form of images with their corresponding class names. Based on these data, the goal is to estimate a VLM’s zero-shot image classification capacity ranking among the VLM zoo on the target dataset.
LOVM encounters two challenges stemming from the inherent heterogeneity in models and datasets. The first challenge is the Modality Gap across different modal features extracted by a VLM. Since the visual and textual features extracted by VLMs tend to cluster into two distinct groups and have gap vectors between them (Liang et al., 2022), using text data as image proxies to rank VLMs is inaccurate. The second challenge is the Capability Gap between the VLM’s overall ranking and its ranking in the specific target dataset. Owing to the VLM’s performance variation across different datasets, the VLM’s average performance on open-source datasets is hard to reflect its performance on a specific target dataset. Thus, selecting a VLM based solely on its general strength proves to be an ineffective strategy.
In this paper, we propose VLM Selection With gAp Bridging (Swab) to address both gaps. The key idea is to reuse the statistics from open-source datasets to estimate the statistics on the target dataset, which mitigates the negative impact of these two gaps. In particular, Swab first uses optimal transport to calculate the transport matrix based on textual similarity between class names of open-source and target datasets. After applying VLMs on open-source datasets to calculate VLMs’ statistics, i.e., the class-specific modality gap vectors and performance rankings of different VLMs, Swab utilizes these statistics to estimate the same type of statistics on the target dataset. After that, Swab uses the estimated gap vectors to align the features of text data with the features of images from the corresponding category, which bridges the modality gap. The estimated VLMs’ ranking also improves the prediction of their rankings on the target task, bridging the capacity gap. The main contributions of our paper are:
  • We analyze two key challenges in LOVM, which are the modality gap across VLM’s different modal features and the capability gap between the VLM’s overall ranking and its ranking on a specific target dataset.
  • We propose Swab, which utilizes optimal transport to transform useful statistics of VLMs on open-source datasets to the target dataset to bridge two gaps.
  • Experimental results on a LOVM benchmark composed of a wide range of VLMs and image classification datasets demonstrate the effectiveness of Swab.
2 Preliminary
[…]
2.1 Selecting VLMs from a Model Zoo
Zero-Shot Image Classification of VLM. Assume there is a pre-trained VLM $f=(f^{I},f^{T})$ consisting of an image encoder $f^{I}$ and a text encoder $f^{T}$. Given an image classification dataset $\mathcal{T}$ with $k_{\mathcal{T}}$ class names $C_{\mathcal{T}}=\{c_{1}^{\mathcal{T}},\cdots,c_{k_{\mathcal{T}}}^{\mathcal{T}}\}$, we input the class names $C_{\mathcal{T}}$ (probably with templates like “A photo of {class}”) into the VLM’s text encoder $f^{T}$ to get the image classifiers $\{\hat{\bm{t}}_{j}\}_{j=1}^{k_{\mathcal{T}}}$. Then, given a test image ${\bm{x}}_{i}$, we use the image encoder $f^{I}$ to extract its feature $\hat{{\bm{x}}}_{i}$. Finally, we predict the label via the cosine similarity between the image feature $\hat{{\bm{x}}}_{i}$ and image classifiers $\{\hat{\bm{t}}_{j}\}_{j=1}^{k_{\mathcal{T}}}$. The class with the highest cosine similarity to the image is selected as the predicted class $\hat{y}_{i}$. Equation 1 and Equation 2 describe this zero-shot image classification process.
$\hat{{\bm{x}}}_{i}=f^{I}({\bm{x}}_{i}),\ \hat{\bm{t}}_{j}=f^{T}(c_{j}^{\mathcal{T}}).$ (1)
$\hat{y}_{i}=f(\bm{x}_{i},C_{\mathcal{T}})=\mathop{\rm argmax}_{c_{j}^{\mathcal{T}}\in[C_{\mathcal{T}}]}\frac{\hat{{\bm{x}}}_{i}^{\top}\hat{\bm{t}}_{j}}{\|\hat{{\bm{x}}}_{i}\|\cdot\|\hat{\bm{t}}_{j}\|}.$ (2)
VLM Zoo. In recent years, there have emerged a large number of (pre-trained) VLMs. Assume a collection of $M$ VLMs constitute a VLM Zoo $\mathcal{M}$:
=$\mathcal{M}=\left\{f_{m}=\left(f_{m}^{I},f_{m}^{T}\right)\right\}_{m=1}^{M}.$ (3)
The capability of $f_{m}$ is determined by three key factors: the model architecture (e.g., Transformer (Vaswani et al., 2017), ConvNeXt (Liu et al., 2022)), the pre-trained dataset (e.g., LAION-400M (Schuhmann et al., 2021), MS-COCO (Lin et al., 2014)), and the training method (e.g., contrastive loss (Radford et al., 2021), caption loss (Yu et al., 2022)). Combinations of these factors result in “good and diverse” VLMs in $\mathcal{M}$. Given a dataset $\mathcal{T}$, it is probable to find a suitable VLM from the VLM zoo with high zero-shot image classification performance on $\mathcal{T}$.
Language-Only VLM Selection (LOVM). Rather than using the images from the target dataset, LOVM focuses on the zero-shot scenario where only the text data such as class names $C_{\mathcal{T}}$ from the target dataset are available for VLM selection. Besides, we can obtain some open-source image classification datasets $\mathcal{S}$. The set of class names in $\mathcal{S}$ is $C_{\mathcal{S}}=\{c_{1}^{\mathcal{S}},\cdots,c_{k_{\mathcal{S}}}^{\mathcal{S}}\}$, and the $D_{\mathcal{S}}^{I}$ denote the labelled images in these classes. Given a target task $\mathcal{T}$, the VLM selection method $h$ estimates the zero-shot classification ability of $f_{m}$ based on $C_{\mathcal{T}}$, $C_{\mathcal{S}}$, and $D_{\mathcal{S}}^{I}$ as in Equation 4. Here $m\in[1,\cdots,M]$.
$\hat{r}_{m,\mathcal{T}}=h(f_{m}\;|\;C_{\mathcal{T}},C_{\mathcal{S}},D_{\mathcal{S}}^{I}).$ (4)
$\hat{r}_{m,\mathcal{T}}$ is the predicted ranking of the $m$-th VLM $f_{m}$ on the target dataset $\mathcal{T}$. The higher the ranking, the more probable $f_{m}$ achieves higher zero-shot image classification performance on the target dataset $\mathcal{T}$.
Given the test image set $D_{\mathcal{T}}^{I}$ of the target dataset $\mathcal{T}$ with $|D_{\mathcal{T}}^{I}|$ images, the zero-shot image classification accuracy $p_{m,\mathcal{T}}$ of $f_{m}$ is calculated by:
$p_{m,\mathcal{T}}$ $=\frac{1}{|D_{\mathcal{T}}^{I}|}\underset{(\bm{x}_{i},y_{i})\in D_{\mathcal{T}}^{I}}{\sum}\mathbb{I}\left(y_{i}=f_{m}\left(\bm{x}_{i},C_{\mathcal{T}}\right)\right).$ (5)
$f_{m}\left(\bm{x}_{i},C_{\mathcal{T}}\right)$ represents the predicted class with the same manner as Equation 1 and Equation 2. $\mathbb{I}(\cdot)$ is the indicator function, which outputs 1 if the condition is satisfied, and 0 otherwise. Based on $\{p_{m,\mathcal{T}}\}_{m=1}^{M}$, we obtain the true ranking of $M$ VLMs $\bm{r}_{\mathcal{T}}=[r_{1,\mathcal{T}},\ldots,r_{M,\mathcal{T}}]$ by assigning higher ranking $r$ to models with higher accuracy $p$. Since in practical applications, we can’t obtain the test images set $D_{\mathcal{T}}^{I}$ in advance, the goal of LOVM is to make the predicted ranking $\hat{\bm{r}}=[\hat{r}_{1,\mathcal{T}},\cdots,\hat{r}_{M,\mathcal{T}}]$ be an accurate estimation of the ground truth ranking $\bm{r}_{\mathcal{T}}$ so that the best VLM can be selected from the VLM zoo.
Evaluation of LOVM Methods. We measure the performance of the LOVM algorithm by comparing the ranking similarity between $\bm{r}_{\mathcal{T}}$ and $\hat{\bm{r}}_{\mathcal{T}}$. Specifically, we calculate the Top-5 Recall $R_{5}$ (ranges from 0 to 1) and Kendall’s Rank Correlation $\tau$ (ranges from -1 to 1). The larger these two metrics are, the better the LOVM algorithm is.
2.2 Possible Paradigms for LOVM
Non-Learning-based LOVM. There are three main paradigms for LOVM. The first paradigm is to neglect the visual encoder and select VLM solely on texts. In detail, we can utilize ChatGPT (Ouyang et al., 2022) to generate auxiliary texts $\tilde{D}_{\mathcal{T}}$ based on class names $C_{\mathcal{T}}$ of $\mathcal{T}$. For example, a generated text for the class “lion” is “A lion is sleeping”. More details are described in the Appendix subsection B.1. These class-specific texts act as substitutes for images, which are referred to as “image proxies” in the following text. Then, whether a VLM $f_{m}$ fits $\mathcal{T}$ could be measured by transferability metrics, e.g., H-Score (Bao et al., 2019) and LogME (You et al., 2021), between the VLM’s text encoder $f_{m}^{T}$ and generated text dataset $\tilde{D}_{\mathcal{T}}$. The second solution relies on the general performance of a certain VLM $f_{m}$. We use some open-source datasets to measure a VLM’s general performance. If $f_{m}$ achieves high zero-shot classification performance over open-source datasets, then it is expected to be competitive on $\mathcal{T}$. The latent assumption is that a VLM’s ranking is relatively consistent across tasks.
Learning-based LOVM. In addition, the ability of a VLM could also be predicted based on a ranker model $f_{R}$. The input of $f_{R}$ is a vector $\bm{s}_{m,\mathcal{T}}$, depicting the dataset-specific representation of $f_{m}$ on $\mathcal{T}$, while the output of $f_{R}$ is the relative/absolute performance $\hat{p}_{m,\mathcal{T}}\in\mathbb{R}$ of $f_{m}$ on $\mathcal{T}$. The $f_{R}$ could be learned on open-source datasets $\mathcal{S}$ (Zhang et al., 2023; Zohar et al., 2023). Due to the availability of both class names $C_{\mathcal{S}}$ and images $D_{\mathcal{S}}^{I}$ in the open-source dataset $\mathcal{S}$ such as ImageNet (Deng et al., 2009), we can calculate each VLM’s representation $\{\bm{s}_{m,n}\}_{m=1,n=1}^{m=M,n=N}$ and true zero-shot image classification accuracy $\{p_{m,n}\}_{m=1,n=1}^{m=M,n=N}$. Here $N$ refers to the number of datasets in $\mathcal{S}$. After constructing the train set, the ranker model $f_{R}$ is learned based on the $\{\bm{s}_{m,n},p_{m,n}\}_{m=1,n=1}^{m=M,n=N}$:
$\min_{f_{R}}\;\sum_{m=1}^{M}\sum_{n=1}^{N}\;\ell(f_{R}(\bm{s}_{m,n}),p_{m,n}).$ (6)
$\ell$ is a loss function that measures the discrepancy between the prediction and the ground-truth. Given $\mathcal{T}$, the learned $f_{R}$ is able to predict the performance $\{\hat{p}_{m,\mathcal{T}}\}_{m=1}^{M}$ over $\{\bm{s}_{m,\mathcal{T}}\}_{m=1}^{M}$ via $\hat{p}_{m,\mathcal{T}}=f_{R}(\bm{s}_{m,\mathcal{T}})$. Finally, we can get the predicted VLMs’ ranking $\hat{\bm{r}}$ based on $\{\hat{p}_{m,\mathcal{T}}\}_{m=1}^{M}$. The representation $\bm{s}_{m,\mathcal{T}}$ is one of the keys in this paradigm, and ModelGPT (Zohar et al., 2023) calculates values $\bm{s}_{m,\mathcal{T}}$ via the capability of a VLM’s text encoder $f_{m}^{T}$.
ModelGPT uses generated text data $\tilde{D}_{\mathcal{T}}$ for $\mathcal{T}$ as substitutes for images to calculate some metrics, which measures the zero-shot ability of $f_{m}$ on unseen images by the classification ability of $f_{m}$ on $\tilde{D}_{\mathcal{T}}$:
$s_{m,\mathcal{T}}^{i}={\rm Metric}_{i}\left(f_{m},\tilde{D}_{\mathcal{T}}\right).$ (7)
Here $\text{Metric}_{i}$ indicates the $i$-th metrics function such as Top-1 Accuracy and F1-Score. For example, the Top-1 Accuracy $s_{m,\mathcal{T}}^{1}$ could be calculated in a similar manner as Equation 2, with the only difference being that the features for the $i$-th text $t_{i}$ in $\tilde{D}_{\mathcal{T}}$ are extracted using a text encoder $f_{m}^{T}$:
$s_{m,\mathcal{T}}^{1}=\frac{1}{|\tilde{D}_{\mathcal{T}}|}\underset{(t_{i},y_{i})\in\tilde{D}_{\mathcal{T}}}{\sum}\mathbb{I}\left(y_{i}=f_{m}(t_{i},C_{\mathcal{T}})\right).$ (8)
Besides, ModelGPT uses some metrics for assessing the features’ quality extracted by the VLM’s text encoder $f_{m}^{T}$. More details are in the Appendix subsection B.2. Moreover, the zero-shot classification performance of $f_{m}$ on open-source datasets $\mathcal{S}$ is also included in $\bm{s}_{m,\mathcal{T}}$ as a general ability measure of $f_{m}$. ModelGPT implements $f_{R}$ as a linear model and sets $\ell$ as the square loss.
Figure 2: Validation Experiments on the Modality Gap and Capacity Gap. (a) Predicted VLMs’ zero-shot image classification accuracy based on auxiliary text data vs. VLM’s true accuracy is based on test images. Each point in the graph represents a model. We calculate Kendall’s tau correlation coefficient $\tau$ and the Mean Absolute Error (MAE) between the predicted accuracy and true accuracy. From the result, we can find that the predicted accuracy calculated by using auxiliary text data poorly aligns with the true accuracy, indicating these text data are ineffective image proxies. (b) We calculate the zero-shot image classification performance rankings of 35 VLMs across 23 datasets. $\bm{r_{i}}$ refers to the vector composed of the $i$-th VLM’s ranking across 23 datasets. We compute the average standard deviations and the mean differences between maximum and minimum values for these VLM’s ranking vectors. The result shows the performance of a VLM varies greatly across different datasets.
2.3 Analysis of the Two Gaps in LOVM
There are two main challenges that limit the application of the aforementioned paradigms in LOVM. The first one is the Modality Gap across different modalities’ features in VLM, and the second is the Capacity Gap between VLM’s average performance and dataset-specific performance.
Modality Gap. As described in subsection 2.2, some methods like H-Score, LogME, and ModelGPT utilize the ChatGPT generated auxiliary texts $\tilde{D}_{\mathcal{T}}$ as image proxies to calculate metrics that measure the zero-shot accuracy on the target dataset $\mathcal{T}$. In other words, the zero-shot classification ability across text and image modalities is estimated by the intra-modality classification ability. The latent assumption is that the generated texts and their corresponding images are closely aligned in VLM’s feature space. However, this assumption is difficult to meet (Liang et al., 2022), and VLM’s features are more likely to cluster according to their modalities. In particular, we define the modality gap vector $\bm{g}$ between the features of an image-text pair $({\bm{x}}_{i},t_{i})$ as:
$\bm{g}_{m,i}\coloneqq f_{m}^{I}({\bm{x}}_{i})-f_{m}^{T}(t_{i})\;.$ (9)
Values in the gap vector are far from zero in most cases. We name this phenomenon as Modality Gap in LOVM, which makes the scores on $\tilde{D}_{\mathcal{T}}$ hard to reveal the true zero-shot image classification capability of a VLM on a given dataset.
We conduct a validation experiment on ImageNet with 35 VLMs provided by (Zohar et al., 2023). For each VLM, we first generate 50 auxiliary texts per class as $\tilde{D}_{\mathcal{T}}$ and then calculate the predicted Top-1 accuracy via Equation 8. Next, we use test images to calculate the VLM’s true Top-1 accuracy. The consistency between the predicted Top-1 accuracy and true zero-shot image classification accuracy $p_{m,\mathcal{T}}$ is measured by the Kendall Rank Correlation ($\tau$, higher is better) and Mean Absolute Error (MAE, lower is better). The results are shown in Figure 2, where each point represents a VLM. It can be observed that the predicted accuracy derived from auxiliary texts $\tilde{D}_{\mathcal{T}}$ does not closely match the true zero-shot accuracy, indicating that these generated auxiliary texts in $\tilde{D}_{\mathcal{T}}$ are not effective proxies for images.
To make the auxiliary texts act as better image proxies, one intuitive idea is to estimate the gap vector $\bm{g}$ for each image-text pair. Given the vector, we can add it to the feature $f_{m}^{T}(t_{i})$ of the text $t_{i}$ to eliminate the modality gap, which may further lead to more accurate scores $s_{m,\mathcal{T}}^{i}$ in Equation 7. However, the gap vector cannot be calculated directly without the images from the target dataset. Furthermore, gap vectors for different classes are diverse, so using a shared vector across all datasets may not be a good choice in LOVM.
Figure 3: The workflow of Swab. Swab first uses optimal transport to construct the transport matrix $\bm{\gamma}^{*}\in\mathbb{R}^{k_{\mathcal{S}}\times k_{\mathcal{T}}}$ based on the textual semantic similarity between the open-source datasets’ classes $C_{\mathcal{S}}=\{c_{1}^{\mathcal{S}},\cdots,c_{k_{\mathcal{S}}}^{\mathcal{S}}\}$ and target dataset’s classes $C_{\mathcal{T}}=\{c_{1}^{\mathcal{T}},\cdots,c_{k_{\mathcal{T}}}^{\mathcal{T}}\}$. Then Swab uses the transport matrix to estimate the VLM’s class-specific gap vectors $\{\bm{g}_{m,1},\cdots\}$ on the target dataset $\mathcal{T}$ based on VLM’s gap vectors $\bm{G}_{m}^{\mathcal{S}}\in\mathbb{R}^{k_{\mathcal{S}}\times d}$ on the open-source datasets. The estimated gap vectors are used to modify text samples for acting as better image proxies, which are input to the Ranker Model $f_{R}$ to predict the $m$-th VLM’s performance $\hat{p}_{m,\mathcal{T}}^{1}$ on the target dataset. Besides, Swab also predicts VLM’s performance based on the VLM’s class-specific ranking $\bm{r}_{m}^{\mathcal{S}}\in\mathbb{R}^{k_{\mathcal{S}}}$ on open-source datasets using the transport matrix $\bm{\gamma}^{*}$. At last, Swab ensembles these predicted performance $\hat{p}_{m,\mathcal{T}}^{1}$ and $\hat{p}_{m,\mathcal{T}}^{2}$ to get the VLM’s final ranking prediction.
Capacity Gap. To select one VLM from the model zoo given a target dataset, one direct approach is to select the “strongest” VLM in all cases. For example, we may first estimate the VLM’s zero-shot classification ability on open-source datasets and then utilize the VLM with the highest performance, which is described in subsection 2.2. Will a VLM’s average ranking on the open-source datasets reveal its true ranking on the target dataset? Our empirical analyses indicate that there exists a discrepancy between the VLM’s overall ranking and its ranking on a specific target dataset. We name the difference between the average ability and the specific ability as the Capacity Gap, which results from the fact that a VLM’s performance fluctuates significantly across various datasets.
To verify the claim, we test 35 VLMs on 23 target datasets provided by (Zohar et al., 2023) and obtain the rankings of each VLM across these datasets. Based on these ranking results, we calculate the average standard deviation and the mean difference between each VLM’s maximum and minimum ranking. The details are illustrated in the right part of Figure 2. The high value of average standard deviation and range difference demonstrate that a VLM’s zero-shot classification ability depends on the property of the target dataset. For example, the mean difference between the maximum and the minimum ranking of one VLM is 28, while the total number of evaluated VLMs is 35. Therefore, the top-performing VLM in one dataset could likely be among the worst in another.
To bridge such a capacity gap, one possible solution is to take account of the ranking of all VLMs on a related dataset. In other words, instead of using the general performance ranking, the specific ranking on some sampled datasets from open-source datasets may be helpful. The main challenge is to figure out which open-source dataset is similar to the target one, and transform the VLM’s ranking on that dataset to the target dataset.
Summary. We emphasize two kinds of gaps in LOVM, i.e., the modality gap across features of different modalities generated by a VLM, and the capability gap between a VLM’s overall ranking and its ranking given a specific target dataset. Both two gaps pose obstacles to some methods, such as LogME and ModelGPT, and degrade their abilities in VLM selection. Moreover, those intuitive approaches to bridge the gaps still face challenges.
3 VLM Selection with Gap Bridging
To mitigate the impact of both gaps on LOVM and integrate non-learning-based and learning-based LOVM methods, we propose VLM Selection With gAp Bridging (Swab). The key idea of Swab lies in bridging the modality and capacity gaps by leveraging useful class-level statistics of VLMs based on open-source datasets. By measuring the textual similarity between the target dataset’s class names and those in open-source datasets, we construct a bridge matrix, which indicates how classes from the open-source datasets can depict a particular VLM’s characteristics for the target dataset. We then estimate the gap vectors between image and text modalities, which rectifies the text-derived scores in ModelGPT. In addition, we assess the VLM’s performance ranking for a specified dataset. Two estimations are merged to achieve a more accurate language-only VLM selection. The workflow of Swab is illustrated in Figure 3.
3.1 Construct the Bridge Using Optimal Transport
Benefiting from the open-source datasets, some useful class-level statistics, such as modality gap vectors and zero-shot classification accuracy of a certain VLM, could be calculated, which may help the ability estimation of a VLM on the target dataset. Instead of reusing the class-level statistics in a uniform manner, our Swab automatically determines the semantic relationship between open-source datasets’ classes and target dataset’s classes with Optimal Transport (Cuturi, 2013; Peyré & Cuturi, 2019).
Recall that the sets of class names of the open-source datasets and the target dataset are $C_{\mathcal{S}}=\{c^{\mathcal{S}}_{i}\}^{k_{\mathcal{S}}}_{i=1}$ and $C_{\mathcal{T}}=\{c^{\mathcal{T}}_{i}\}^{k_{\mathcal{T}}}_{i=1}$, respectively. The semantic relationship between two classes could be measured by the textual similarity between their class names. In detail, we use a pre-trained text encoder $\phi$ (e.g., MPNet (Song et al., 2020)), which extracts same-dimensional features for all previous class names, i.e., $\{\phi(c_{1}^{\mathcal{S}}),\cdots,\phi(c_{k_{\mathcal{S}}}^{\mathcal{S}})\}$ and $\{\phi(c_{1}^{\mathcal{T}}),\cdots,\phi(c_{k_{\mathcal{T}}}^{\mathcal{T}})\}$. Then, the distance between the $i$-th class in the open-source datasets and the $j$-th class in the target dataset could be measured based on their cosine similarity:
${\rm cost}_{ij}=1-\frac{\phi(c_{i}^{\mathcal{S}})^{\top}\phi(c_{j}^{\mathcal{T}})}{\|\phi(c_{i}^{\mathcal{S}})\|\cdot\|\phi(c_{j}^{\mathcal{T}})\|}\;.$ (10)
Elements in ${\rm cost}\in\mathbb{R}_{+}^{k_{\mathcal{S}}\times k_{\mathcal{T}}}$ reveal how much the extracted text embeddings of these two classes are close to each other. In practical implementation, we utilize the exponential of each element in ${\rm cost}$ with the base of the natural constant $e$, which enlarges the difference of its values.
Values in the matrix ${\rm cost}$ indicate the extent of the cost required to reuse the element from a certain open-source class in a specific target class. In other words, if two classes are distant based on their textual semantics, it is more probable that reusing the statistics from that open-source class to help a target class is difficult. We take advantage of Optimal Transport (OT) to reuse more information between semantically similar classes:
$\bm{\gamma}^{*}=\underset{\bm{\gamma}\in\mathbb{R}_{+}^{k_{\mathcal{S}}\times k_{\mathcal{T}}}}{\rm argmin}\sum_{i,j}\gamma_{i,j}\;{\rm cost}_{i,j}$ (11)
[…]
The cost matrix ${\rm cost}$ quantifies the expense of moving elements between all class pairs, and $\bm{\gamma}^{*}\in\mathbb{R}^{k_{\mathcal{S}}\times k_{\mathcal{T}}}$ is the transport matrix. OT minimizes the cost indicated by the matrix ${\rm cost}$ and moves elements from one distribution $\bm{u}$ to another $\bm{v}$. In Swab, we define $\bm{u}$ and $\bm{v}$ as uniformly distributed vector $\bm{u}=\bm{1}/k_{\mathcal{S}}\in\mathbb{R}^{k_{\mathcal{S}}}$ and $\bm{v}=\bm{1}/k_{\mathcal{T}}\in\mathbb{R}^{k_{\mathcal{T}}}$. This indicates that we treat all classes as equally important. We may also incorporate prior knowledge of class importance to define $\bm{u}$ and $\bm{v}$.
The solution $\bm{\gamma}^{*}$ of the OT problem in Equation 11 could be solved efficiently (Flamary et al., 2021), and $\bm{\gamma}^{*}$ acts as a bridge matrix between open-source classes and target classes. Usually, the smaller ${\rm cost}_{i,j}$ is, the larger the corresponding element $\bm{\gamma}^{*}_{i,j}$ obtained by OT, indicating statistics of the $i$-th open-source class may help more when we estimate the statistics of the $j$-th target class.
3.2 Bridge the Modality Gap and Capacity Gap
In Swab, we take advantage of the class relationship indicated by $\bm{\gamma}^{*}_{i,j}$ and bridge two kinds of gaps.
Bridge the Modality Gap. Given the $m$-th VLM $f_{m}$ in the model zoo, we want to estimate the modality gap $\bm{g}_{m,j}$ between the extracted image and text features for the $j$-th class in the target dataset $\mathcal{T}$ to bridge the modality gap. However, there are no images from the target dataset, so we can’t directly calculate the gap vectors using image-text pairs. Swab estimates the target dataset’s gap vectors based on the open-source datasets’ gap vectors with $\bm{\gamma}^{*}_{i,j}$.
Given the $k$-th open-source class $c_{k}^{\mathcal{S}}$, we collect the set of images $D_{\mathcal{S}_{k}}^{I}$ from this class
$D_{\mathcal{S}_{k}}^{I}=\left\{\left(\bm{x}_{i},y_{i}\right)|\left(\bm{x}_{i},y_{i}\right)\in D_{\mathcal{S}}^{I},\ y_{i}=c_{k}^{\mathcal{S}}\right\}.$ (12)
We denote $|D_{\mathcal{S}_{j}}^{I}|$ as the images number in $D_{\mathcal{S}_{j}}^{I}$. Then, the modality gap vector $\bm{g}_{m,k}^{\mathcal{S}}$ for $c_{k}^{\mathcal{S}}$ and $f_{m}$ could be obtained in a similar manner as Equation 9:
$\bm{g}_{m,k}^{\mathcal{S}}=\frac{1}{|D_{\mathcal{S}_{k}}^{I}|}\underset{(\bm{x}_{i},y_{i})\in D_{\mathcal{S}_{k}}^{I}}{\sum}\left(\frac{f_{m}^{I}\left(\bm{x}_{i}\right)}{\|f_{m}^{I}\left(\bm{x}_{i}\right)\|}-\frac{f_{m}^{T}\left(c_{k}^{\mathcal{S}}\right)}{\|f_{m}^{T}\left(c_{k}^{\mathcal{S}}\right)\|}\right)\;.$ (13)
$\bm{g}_{m,k}^{\mathcal{S}}$ is the average value of the difference between the normalized class name embedding and all normalized image embeddings from the class $c_{k}^{\mathcal{S}}$. Then, the gap vectors of all open-source classes $\{\bm{g}_{m,1}^{\mathcal{S}},\cdots,\bm{g}_{m,k_{\mathcal{S}}}^{\mathcal{S}}\}$ can be obtained given $f_{m}$. We use a matrix $\bm{G}_{m}^{\mathcal{S}}\in\mathbb{R}^{k_{\mathcal{S}}\times d_{m}}$ to represent those $k_{\mathcal{S}}$ gap vectors for the $m$-th VLM’s, and $d_{m}$ is the dimensionality of features extracted by $f_{m}$.
The gap vectors $\{\bm{g}_{m,1},\cdots,\bm{g}_{m,k_{\mathcal{T}}}\}$ for the target dataset could be estimated based on the $\bm{G}_{m}^{\mathcal{S}}$ and the class-wise relationship. If two classes are semantically similar, then we can reuse the gap vector from the similar class. We set the gap vector for the $j$-th target class $\bm{g}_{m,j}$ as a weighted sum of $\bm{G}_{m}^{\mathcal{S}}$, and the weight comes from $\bm{\gamma}^{*}_{i,j}$:
$\bm{g}_{m,j}=|C_{\mathcal{T}}|\;{\bm{G}^{\mathcal{S}}_{m}}^{\top}\bm{\gamma}^{*}_{:,j}.$ (14)
$\bm{\gamma}^{*}_{:,j}$ is the $j$-th column of $\bm{\gamma}^{*}$, whose elements are the relationship between the $j$-th target class to all open-source classes. We also use scaling factors $|C_{\mathcal{T}}|$ to ensure that for each target class, the sum of $\bm{\gamma}^{*}_{:,j}$ equals 1. This scale operation has also been used in previous work (Ye et al., 2018, 2021).
After that, we modify the essential step of ModelGPT in Equation 7, where the metrics over the generated auxiliary texts $\tilde{D}_{\mathcal{T}}$ are measured. We add the gap vector $\bm{g}_{m,j}$ to the embeddings of the auxiliary texts $\tilde{D}_{\mathcal{T}}^{j}$ from the $j$-th class in the target dataset:
$\tilde{\bm{t}}_{m,i}=f^{T}_{m}(t_{i})+\bm{g}_{m,j}\;,\;\;\forall t_{i}\in\tilde{D}_{\mathcal{T}}^{j}\;.$ (15)
The modified text embedding $\tilde{\bm{t}}_{m,i}$ serves as better image proxies. In other words, classification metrics on $f^{T}_{m}(t_{i})$ only reveal the discerning ability of the text encoder of $f_{m}$, which is far from the (cross-modal) zero-shot classification ability due to the modality gap. By bridging such a gap with modified text embedding, classification metrics on $\tilde{\bm{t}}_{m,i}$ are closer to the classification metrics on images with textual classifier. Therefore, we use $\{\tilde{\bm{t}}_{m,1},\cdots\}$ in Equation 15 as better inputs to calculate $\bm{s}_{m,\mathcal{T}}$ in Equation 7. The updated score vectors $\bm{s}_{m,\mathcal{T}}$ are then input to the ranker model $f_{R}$, which is able to get more accurate VLM’s performance prediction $\hat{p}_{m,\mathcal{T}}$. Based on the performance prediction $\{\hat{p}_{m,\mathcal{T}}\}_{m=1}^{M}$, we can obtain the VLMs’ ranking:
$\bm{\hat{r}}_{k,\mathcal{T}}$ $=[\hat{r}_{1,\mathcal{T}},\cdots,\hat{r}_{M,\mathcal{T}}]$ (16)
$=\text{Ranking}\left(\left[\hat{p}_{1,\mathcal{T}},\cdots,p_{M,\mathcal{T}}\right]\right)\;.$
$\text{Ranking}(\cdot)$ transforms the accuracy values into their performance ranking within the VLM Zoo. $\bm{r}_{k}^{\mathcal{S}}\in\mathbb{Z}_{+}^{M}$ contains integer values from 1 to $M$.
Bridge the Capacity Gap. Whether the $m$-th VLM $f_{m}$ fits the target task $\mathcal{T}$ could also be determined by the performance of $f_{m}$ on the open-source datasets related to $\mathcal{T}$. As we mentioned in subsection 2.3, the performance of a VLM varies from one dataset to another, so we calculate the VLM’s class-specific performance ranking over the whole open-source datasets. Given the $k$-th open-source class $c_{k}^{\mathcal{S}}$ and the corresponding set of images $D_{\mathcal{S}_{k}}^{I}$ as in Equation 12, we calculate the zero-shot classification accuracy $p_{m,k}^{\mathcal{S}}$ via Equation 2. Based on the performance rankings of all $M$ VLMs, we obtain each VLMs’ ranking on the $k$-th open-source class $c_{k}^{\mathcal{S}}$ using Equation 17. We calculate VLMs’ ranking for each of the $k_{\mathcal{S}}$ open-source classes similarly.
$\bm{r}_{k}^{\mathcal{S}}$ $=[r_{1,k}^{\mathcal{S}},\cdots,r_{M,k}^{\mathcal{S}}]$ (17)
$=\text{Ranking}\left(\left[p_{1,k}^{\mathcal{S}},\cdots,p_{M,k}^{\mathcal{S}}\right]\right)\;.$
Next, we consider estimating the ranking of a certain VLM $f_{m}$ on the target dataset $\mathcal{T}$. By re-organizing the ranking values in Equation 17, the performance ranking vector of $f_{m}$ on all $k_{\mathcal{S}}$ open-source classes is $\bm{r}_{m}^{\mathcal{S}}=[r_{m,1}^{\mathcal{S}},\cdots,r_{m,k_{\mathcal{S}}}^{\mathcal{S}}]\in\mathbb{Z}_{+}^{k_{\mathcal{S}}}$. If $f_{m}$ has a higher ranking on certain open-source classes and those classes are related to the classes in the target set, the performance ranking of $f_{m}$ on the target dataset may also be high.
Thus, we perform a weighted sum of ranking values in $\bm{r}_{m}^{\mathcal{S}}$ and assign larger weights to those classes related to the target dataset. Since the semantic similarity between open-source and target classes is measured by $\bm{\gamma}^{*}$, we apply
$\hat{\bm{r}}_{m,\mathcal{T}}=\bm{r}_{m}^{\mathcal{S}}\;\bm{\gamma}^{*}\;.$ (18)
Elements in $\hat{\bm{r}}_{m,\mathcal{T}}\in\mathbb{R}^{k_{\mathcal{T}}}$ are the predicted ranking of the m-th VLM $f_{m}$ for $k_{\mathcal{T}}$ target classes. Since we only need the relative order of ranks, there is no additional scale factor in Equation 18. After that, we average class-specific ranking values in $\hat{\bm{r}}_{m,\mathcal{T}}$, and use the mean of $\hat{\bm{r}}_{m,\mathcal{T}}$ to denote the ability of $f_{m}$ on the target dataset $\mathcal{T}$. In summary, we take account of the VLM ranks on related datasets, and such class-specific ranking estimation bridges the capacity gap in VLM selection.
3.3 Summary of Swab
As is described in section 3.2, given a target dataset $\mathcal{T}$ and a VLM $f_{m}$, we denote the two performance ranking predictions by bridging the modality gap and capacity gap as $\hat{r}^{1}_{m,\mathcal{T}}$ and $\hat{r}^{2}_{m,\mathcal{T}}$, respectively. $\hat{r}^{1}_{m,\mathcal{T}}$ is predicted based on ModelGPT with our modified embeddings of the generated auxiliary texts for $\mathcal{T}$. $\hat{r}^{2}_{m,\mathcal{T}}$ is predicted by the weighted sum of VLM’s class-specific ranking values on the open-source datasets. These two predictions respectively originate from non-learning-based and learning-based LOVM methods. We ensemble two predictions together and achieve a more accurate model ranking estimation. We utilize the weighted Borda Count to aggregate two rankings:
$\hat{r}^{\rm ens}_{m,\mathcal{T}}=\alpha\cdot\hat{r}^{1}_{m,\mathcal{T}}+\left(1-\alpha\right)\cdot\hat{r}^{2}_{m,\mathcal{T}}.$ (19)
Ultimately, Swab determines the final predicted ranking of the VLMs in the VLMs Zoo based on $\hat{r}^{\rm ens}_{m,\mathcal{T}}$. The pseudo-code of Swab is listed in Algorithm 1.
4 Experiments
We evaluate Swab on LOVM benchmark (Zohar et al., 2023) and analyze whether bridging the modality gap and capacity gap helps VLM selection.
4.1 Evaluation on LOVM Benchmark
Setups. We follow LOVM (Zohar et al., 2023) to use a VLM Zoo with 35 pre-trained VLMs, which differ in aspects such as model architecture, pre-training datasets and training methods. We evaluate different methods on 23 datasets, i.e. ImageNet (Deng et al., 2009), Aircraft (Maji et al., 2013), CIFAR100 (Krizhevsky & Hinton, 2009) and so on. We obtain VLM’s ground truth ranking based on VLM’s Top-1 Accuracy calculated on target dataset’s test image set.
Baseline. We select representative methods for each of the three paradigms mentioned in Section 2.2 as our baselines. For the first paradigm, we use four classic model selection methods including H-Score (Bao et al., 2019), NCE (Tran et al., 2019), LEEP (Tran et al., 2020) and LogME (You et al., 2021). For the second paradigm, we use the ranking of VLM on ImageNet (INB) and the average ranking (Average Rank) on all datasets except the current target dataset in the LOVM Benchmark as our baselines, respectively. For the third paradigm, we compare our method with ModelGPT (Zohar et al., 2023).
Evaluations. We use Top-5 Recall and Kendall’s Rank Correlation to measure the ranking similarity between the predicted model rankings and the ground truth ones to evaluate the VLM selection method’s performance. We also calculate the sum of these two metrics to consider the method’s comprehensive capability.
Implementation Details. For a fair comparison, Swab follow ModelGPT (Zohar et al., 2023) to sequentially extract a target dataset from each of the 23 datasets in the LOVM Benchmark and treat the remaining datasets as open-source datasets. Besides, Swab adopts the approach outlined in ModelGPT to add Gaussian noise to corrupt the embeddings of the target dataset’s generated text data, which can improve the VLM selection method’s performance. Adding noise to features to boost the method’s performance has been used in prior work (Ye et al., 2017). For H-Score, NCE, LEEP, LogME, INB, and Average Rank, we follow the practices of previous work and do not add noise to the model’s inputs. To avoid the randomness introduced by the noise affecting the reliability of the experiment results, we conduct ten repeated experiments using random seeds from 1 to 10 and report the mean performance and standard deviation of different methods in Table 1.
Table 1: Results on LOVM Benchmark. We evaluate our method’s performance over 23 datasets and 35 pre-trained VLMs. The results are averaged over all datasets. Our Swab achieves the best results across all metrics. For methods that involve adding random noise to data features, we report the standard deviation of metrics across 10 experiments to mitigate the impact of randomness on result reliability.
Method $R_{5}(\uparrow)$ $\tau(\uparrow)$ $R_{5}$ + $\tau(\uparrow)$
H-Score 0.174 0.035 0.209
NCE 0.252 -0.029 0.223
LEEP 0.174 0.035 0.209
LogME 0.209 -0.029 0.180
INB 0.452 0.177 0.629
Average Rank 0.452 0.133 0.585
ModelGPT $\text{0.457}\scriptstyle\pm\text{0.009}$ $\text{0.203}\scriptstyle\pm\text{0.018}$ $\text{0.660}\scriptstyle\pm\text{0.015}$
Swab $\textbf{0.486}\scriptstyle\pm\text{0.010}$ $\textbf{0.268}\scriptstyle\pm\text{0.011}$ $\textbf{0.754}\scriptstyle\pm\text{0.003}$
Results Analysis. From Table 1, we can draw the following conclusions: (1) Metric-based non-learning model selection methods such as LogME show poor performance on the LOVM Benchmark. This is primarily because such algorithms rely on the target dataset’s images, thus the modality gap has a greater negative impact on them when using generated text data as a substitute for images. (2) Using open-source datasets is helpful for LOVM. We find that using open-source datasets in a non-learning way (e.g. INB, Average Rank) or a learning way (e.g. ModelGPT) all helps LOVM, since their performance significantly surpasses that of methods not utilizing open-source datasets (e.g. LogME). (3) Despite leveraging more open-source datasets, the performance of Average Rank is worse than INB. This confirms our analysis of the Capacity Gap, which suggests a discrepancy between the average ranking of a VLM and its ranking on a specific dataset. (4) Our Swab achieves the best performance across all evaluation metrics. Notably, our final performance of $R_{5}+\tau$ (0.754) represents a significant improvement of 9.4% over the original SoTA (State-of-The-Art) method ModelGPT (0.660).
4.2 Ablation Study
We conduct ablation studies to demonstrate that bridging the Modality Gap and Capacity Gap are both essential for Swab. Table 2 presents our experiment results, from which we can observe that Swab achieves the best performance when both gaps are bridged simultaneously. The ablation study confirms our analysis.
Table 2: Ablation Study of LOVM. Swab-C, Swab-M, and Swab indicates only bridging the Modality Gap, only bridging the Capacity Gap, and bridging both gaps in Swab.
Method $R_{5}(\uparrow)$ $\tau(\uparrow)$ $R_{5}$ + $\tau(\uparrow)$
Swab-C $\text{0.478}\scriptstyle\pm\text{0.012}$ $\text{0.213}\scriptstyle\pm\text{0.011}$ $\text{0.691}\scriptstyle\pm\text{0.007}$
Swab-M $\text{0.459}\scriptstyle\pm\text{0.010}$ $\text{0.251}\scriptstyle\pm\text{0.001}$ $\text{0.710}\scriptstyle\pm\text{0.011}$
Swab $\textbf{0.486}\scriptstyle\pm\text{0.010}$ $\textbf{0.268}\scriptstyle\pm\text{0.011}$ $\textbf{0.754}\scriptstyle\pm\text{0.003}$
4.3 Influence of Key Components in Swab
Will Bridge the Capacity Gap Be Beneficial for VLM Selection? We compare the LOVM performance directly using the VLM’s average ranking on each class of open-source datasets and weighted-sum ranking based on transport matrix $\bm{\gamma}^{*}$. The results are shown in Table 3. From the results, we can find that utilizing class relevance to bridge the Capacity Gap is beneficial for VLM’s Model Selection.
Table 3: Compare results of $\hat{r}^{2}_{m,\mathcal{T}}$ on the LOVM Benchmark before and after bridging the Capacity Gap.
Method $R_{5}(\uparrow)$ $\tau(\uparrow)$ $R_{5}$ + $\tau(\uparrow)$
Average Rank 0.452 0.133 0.585
OT Weighted Rank 0.443 0.275 0.718
Will Bridge the Modality Gap Be Beneficial for VLM Selection? To eliminate the interference of other factors, we solely utilize the learning-based predicted rankings $\hat{r}^{1}_{m,\mathcal{T}}$ in Swab, and the input to the ranker model $f_{m}$ consists only of metrics calculated on the generated text data $\tilde{D}_{\mathcal{T}}$. In this way, the performance of the method depends solely on the quality of the generated text data $\tilde{D}_{\mathcal{T}}$. From the Table 4, we can find that the generated text data $\tilde{D}_{\mathcal{T}}$ become better image proxies after bridging the Modality Gap.
Table 4: Compare results of $\hat{r}^{2}_{m,\mathcal{T}}$ on the LOVM Benchmark before and after bridging the Modality Gap (MG).
Method $R_{5}(\uparrow)$ $\tau(\uparrow)$ $R_{5}$ + $\tau(\uparrow)$
Before Bridging MG 0.249 0.064 0.313
After Bridging MG 0.350 0.196 0.546
Which Kind of Gap Vectors Should We Use? When utilizing the gap vectors from open-source datasets, we have two options: (1) Use the dataset-level mean gap vector calculated on the whole dataset’s image-text pairs. (2) Use the class-level mean gap vector calculated on the corresponding class’s image-text pairs. We hope that the gap vectors used to calculate the mean gap vector are as close to each other as possible, so that their mean vector can serve as a good substitute for the whole set. Based on this idea, we examine the statistical properties of dataset-level gap vectors set and class-level gap vectors set, respectively. We calculate three metrics which include: (1) the standard deviation of these gap vectors’ magnitude; (2) the mean cosine similarity between these gap vectors and their corresponding mean gap vectors; and (3) the standard deviation of these cosine similarities. These metrics reflect the consistency of the gap vectors. Table 5 shows the results.
Table 5: Results of metrics measuring gap vectors’ consistency. M: Magnitude, D: Direction. Metrics are averaged across datasets and classes, respectively.
Gap Vector ImageNet
[…]
Dataset Mean 0.04 0.68 0.07
Class Mean 0.03 0.85 0.05
From the Table 5, we can find that the class-level gap vectors tend to be more consistent, which inspires us to use the class-level mean gap vectors. We also compare the results of $\{\hat{r}^{1}_{m,\mathcal{T}}\}_{m=1}^{M}$ on LOVM Benchmark using the dataset-level mean gap vectors and the class-level mean gap vectors, respectively. The implementation details are the same as Table 4. Table 6 shows the results, which verifies that using the class-level mean gap vectors is a better choice.
Table 6: Results of $\hat{r}^{1}_{m,\mathcal{T}}$ on the LOVM Benchmark using the dataset-level mean gap vectors and class-level mean gap vectors.
Gap Vector $R_{5}(\uparrow)$ $\tau(\uparrow)$ $R_{5}$ + $\tau(\uparrow)$
Dataset Mean 0.338 0.091 0.429
Class Mean 0.350 0.196 0.546
Figure 4: The Spider Charts of a VLM. The Spider Chart measures the model’s capabilities across different domains. The axis represents the model’s average ranking on datasets in different domains. The greater the overlap between the predicted VLM’s spider chart and the actual VLM’s spider chart, the better the model selection method is. Our Swab performs better than ModelGPT.
4.4 Visualizing Swab’s Results by Spider Chart
We select some datasets from six different domains out of the 23 datasets in LOVM Benchmark (see Appendix C.4 for more details). We take the average of the model’s rankings across multiple datasets within the same domain as the model’s ranking in this domain. Afterward, we use ModelGPT and Swab, respectively, to predict the VLM rankings on datasets across different domains, thereby obtaining the predicted rankings of the VLM in six domains. We use a spider chart to visualize the gap between the VLM’s predicted rankings and actual rankings. The different vertices of the spider chart represent the VLM’s ranking in different domains. The score at each vertex indicates the model’s specific ranking. The higher the ranking (closer to 1), the better the model’s performance in that category. Figure 4 displays the spider chart for OpenAI’s CLIP RN50×64. By comparing the VLM’s predicted rankings generated by Swab and ModelGPT with the actual VLM rankings, we find that Swab is more capable of accurately estimating a VLM’s abilities compared to ModelGPT, and the abilities estimated by Swab are very close to the model’s true capabilities.
5 Related Work
Vision-Language Models. Vision-Language Models represent a class of multimodal models adept at correlating textual and visual information. VLMs are pre-trained on extensive text-image pairs using contrastive loss, endowing them with powerful text-image matching capability. Prominent VLMs include CLIP (Radford et al., 2021), ALIGN (Jia et al., 2021), FLAVA (Singh et al., 2022), Florence (Yuan et al., 2021), and GLIP (Li et al., 2022). VLMs possess robust zero-shot image classification capabilities (Radford et al., 2021), which enables its widespread application in tasks characterized by long-tail distributions or those where collecting substantial training data is challenging, such as medical image analysis. Some work (Menon & Vondrick, 2023; Yi et al., 2024) has also shown that by incorporating external knowledge, the zero-shot capabilities of CLIP can be further enhanced. In recent years, the number of open-source VLMs has been increasing (Ilharco et al., 2021). Previous work (Zohar et al., 2023) has pointed out that different VLMs possess varying image classification capabilities. This indicates that the performance of VLMs can vary significantly across different tasks and domains. These models with diverse capabilities constitute a VLM Zoo rich in knowledge. This VLM Zoo enables us to utilize different VLMs for various classification tasks, thereby changing the paradigm of using a single VLM to complete diverse classification tasks. This paper focuses on selecting the most suitable VLM for the target task from the VLM Zoo.
Model Selection.  Previous model selection methods typically estimate a PTM’s performance on a target task by estimating the correlation between the features extracted by the pre-trained model itself (Tran et al., 2019, 2020; You et al., 2021, 2022; Ding et al., 2022; Huang et al., 2022) or a general model (Zhang et al., 2023) and target labels on the target dataset’s samples. However, VLMs are typically used in zero-shot or few-shot scenarios, where target data is limited, making traditional model selection approaches unsuitable for VLMs. Additionally, previous methods have mainly focused on single-modal models, overlooking the characteristics of VLMs. Therefore, in this paper, we concentrate on designing model selection algorithms that are suitable for scenarios with limited data and take into account the characteristics of VLMs.
6 Conclusion
We analyze and address two key challenges in Language-Only VLM Selection (LOVM), which are VLM’s modality gap across different modal features and VLM’s capacity gap between its overall and dataset-specific rankings. Our key insight is that we can reuse the model’s useful statistics on open-source tasks to help the model selection on the target task. Swab utilizes a transport matrix between classes of the target task and open-source datasets to transfer VLM’s class-specific modality gap vectors and class-specific rank from open-source tasks to the target task, which mitigates the negative impacts of these two gaps. Experiment results on the LOVM benchmark show the superiority of our method.
Appendix A LOVM Benchmark Details
LOVM Benchmark (Zohar et al., 2023) consists of 35 pre-trained VLMs and 23 datasets, with a total of $35\times 23=805$ evaluations. For each evaluation, LOVM provides the VLM’s zero-shot image classification accuracy on the corresponding dataset. Therefore, we can get the ground truth performance ranking of 35 VLMs on the 23 datasets.
A.1 VLMs of LOVM Benchmark
To cover as many types of models as possible, the LOVM Benchmark uses OpenCLIP library (Ilharco et al., 2021) to get diverse VLMs. These VLMs differ from each other in terms of the model architecture (ResNet (He et al., 2016), Transformer (Vaswani et al., 2017), ConvNext (Liu et al., 2022)), the pre-trained dataset (OpenAI’s Data (Radford et al., 2021), LAION 2b (Schuhmann et al., 2021)), the training method (loss function/hyperparameter/data augmentation) and so on. Table 7 displays the relevant information of each VLM. The diversity of these VLMs ensures that the experimental results calculated on them can reflect the performance of the VLM model selection algorithm in real-world situations.
Table 7: The detailed information of 35 models used in the LOVM Benchmark. The table comes from  (Zohar et al., 2023).
ID Model Name Dataset Name
1 RN50 RN50 openai WIT
2 RN101 RN101 openai WIT
3 RN50x4 RN50x4 openai WIT
4 RN50-16 RN50x16 openai WIT
5 RN50x64 RN50x64 openai WIT
6 ViT-B-32 ViT-B/32 laion400m_e31 L400m
7 ViT-B-32 ViT-B/32 laion400m_e32 L400m
8 ViT-B-32-quickgelu ViT-B/32 laion400m_e32 L400m
9 ViT-B-32 ViT-B/32 openai WIT
10 ViT-B-32 ViT-B/32 laion2b_s34b_b79k L2b-b
11 ViT-B-32 ViT-B/32 laion2b_e16 L2b-c
12 ViT-B-16 ViT-B/16 laion400m_e32 L400m
13 ViT-B-16 ViT-B/16 openai WIT
14 ViT-B-16-240 ViT-B/16-240 laion400m_e32 L400m
15 ViT-L-14 ViT-L/14 laion400m_e31 L400m
16 ViT-L-14 ViT-L/14 laion400m_e32 L400m
17 ViT-L-14 ViT-L/14 laion2b_s32b_b82k L2b-b
18 ViT-L-14 ViT-L/14 openai WIT
19 ViT-L-14-336 ViT-L/14-336 openai WIT
20 ViT-G-14 ViT-G/14 laion2b_s12b_b42k L2b-a
21 ViT-G-14 ViT-G/14 laion2b_s34b_b88k L2b-a
22 ViT-H-14 ViT-H/14 laion2b_s32b_b79k L2b-b
23 coca_ViT-B-32 CoCa-ViT-B/32 laion2b_s13b_b90k L2b-c
24 coca_ViT-B-32 CoCa-ViT-B/32 mscoco_finetuned_laion2b_s13b_b90k L2b-c + coco
25 coca_ViT-L-14 CoCa-ViT-L/14 laion2b_s13b_b90k L2b-c
26 coca_ViT-L-14 CoCa-ViT-L/14 mscoco_finetuned_laion2b_s13b_b90k L2b-c + coco
27 convnext_base ConvNEXT-B laion400m_s13b_b51k L400m-c
28 convnext_base_w ConvNEXT-BW laion2b_s13b_b82k L2b-d
29 convnext_base_w ConvNEXT-BW laion2b_s13b_b82k_augreg L2b-e
30 convnext_base_w ConvNEXT-BW laion_aesthetic_s13b_b82k L2b-f
31 convnext_base_w_320 ConvNEXT-BW-320 laion_aesthetic_s13b_b82k L2b-f
32 convnext_base_w_320 ConvNEXT-BW-320 laion_aesthetic_s13b_b82k_augreg L2b-g
33 convnext_large_d ConvNEXT-LD laion2b_s26b_b102k_augreg L2b-h
34 convnext_large_d_320 ConvNEXT-LD-320 laion2b_s29b_b131k_ft L2b-i
35 convnext_large_d_320 ConvNEXT-LD-320 laion2b_s29b_b131k_ft_soup L2b-j
A.2 Datasets of LOVM Benchmark
To cover as wide a distribution of image classification tasks as possible, the LOVM Benchmark collects 23 diverse datasets. These datasets differ from each other in terms of the number of categories, category semantics, image domains, and so on. Table 8 displays the relevant information of each dataset. The diversity of these tasks ensures that the experimental results calculated on them can reflect the performance of the VLM model selection method in real-world situations.
Table 8: The detailed information of 23 tasks used in the LOVM Benchmark.
Dataset Classes Task Domain
Imagenet (Deng et al., 2009) 1000 classification natural image
SUN397 (Xiao et al., 2010) 397 scene und. natural image
Country211 (Radford et al., 2021) 211 geolocation natural image
Stanford Cars (Krause et al., 2013) 196 classification natural image
Flowers102 (Nilsback & Zisserman, 2008) 102 classification natural image
CIFAR100 (Krizhevsky & Hinton, 2009) 100 classification natural image
DTD (Cimpoi et al., 2014) 46 classification textural image
RESISC45 (Cheng et al., 2017) 45 classification satellite images
GTSRB (Stallkamp et al., 2011) 43 classification natural image
Oxford Pets (Parkhi et al., 2012) 37 classification natural image
VOC2007 (Everingham et al., 2007) 20 classification natural image
STL10 (Coates et al., 2011) 10 classification natural image
EuroSAT (Helber et al., 2019) 10 classification satellite images
MNIST (LeCun et al., 2010) 10 classification hand-writing
SVHN (Netzer et al., 2011) 10 OCR natural image
CLEVR-C (Johnson et al., 2017) 8 object counting natural image
CLEVR-D (Johnson et al., 2017) 8 distance est. natural image
FER2013 (Goodfellow et al., 2013) 7 fac. exp. rec. natural image
DMLab (Zhai et al., 2020) 6 distance est. synthetic
Retinopathy (Kaggle & EyePacs, 2015) 5 classification retina scan
KITTI (Geiger et al., 2013) 4 distance est. natural image
PCam (Veeling et al., 2018) 2 classification histopathology
Rendered SST2 (Radford et al., 2021) 2 OCR text image
A.3 Evaluation Metrics of LOVM Benchmark
In LOVM, our aim is to maximize the rank similarity between the prediction of VLMs’ performance $\hat{\bm{r}}_{\mathcal{T}}=\{\hat{r}_{m,\mathcal{T}}\}_{m=1}^{M}$ and VLMs’ ground truth performance $\bm{r}_{\mathcal{T}}=\{r_{m,\mathcal{T}}\}_{m=1}^{M}$ on the target dataset, especially the rank similarity of the top 5 VLMs in $\hat{\bm{r}}_{\mathcal{T}}$ and $\bm{r}_{\mathcal{T}}$. This is because we tend to focus only on whether the best models can be chosen. We use the following metrics to evaluate the rank similarity:
  • Top-5 Recall ($R_{5}$) – Top-5 Recall $R_{5}$ measures the model selection algorithm’s accuracy in identifying the true top five best-performing models within its predicted top five models. The calculation method is shown in Equation 21. Here $\text{IND}(\hat{\bm{r}}_{\mathcal{T}}^{5})$ and $\text{IND}(\bm{r}_{\mathcal{T}}^{5})$ indicates the model indices sets of the top 5 VLMs in $\hat{\bm{r}}_{\mathcal{T}}$ and $\bm{r}_{\mathcal{T}}$, respectively. A Top 5 Recall closer to 1 signifies greater accuracy in the predicted rankings.
    $F=\text{IND}(\hat{\bm{r}}_{\mathcal{T}}^{5})\cap\text{IND}(\bm{r}_{\mathcal{T}}^{5}).$ (20)
    $R_{5}=\frac{|F|}{5}.$ (21)
  • Kendall’s Rank Correlation ($\tau$) – Kendall’s Rank Correlation $\tau$ measures the ranking consistency between two ranking lists. Here we mainly focus on whether models within the intersection of the top 5 VLMs in $\hat{\bm{r}}_{\mathcal{T}}$ and $\bm{r}_{\mathcal{T}}$ have consistent rankings. Equation 23 shows the calculation method. The value of $\tau$ ranges from -1 to 1, indicating ranking correlation from complete inconsistency to complete consistency.
    $q_{ij}=\text{sign}(r_{i,\mathcal{T}}-r_{j,\mathcal{T}})\cdot\text{sign}(\hat{r}_{i,\mathcal{T}}-\hat{r}_{j,\mathcal{T}}),$ (22)
    $\tau=\begin{cases}0,&\text{if }|F|<2,\\ \textstyle\frac{2}{|F|(|F|-1)}\sum_{\begin{subarray}{c}i<j;i,j\in F\end{subarray}}q_{ij},&\text{otherwise}.\end{cases}$ (23)
Appendix B ModelGPT Details
ModelGPT is a method proposed for LOVM (Zohar et al., 2023). In this section, we introduce the metrics that ModelGPT used in Equation 7.
B.1 The Generation Process of Auxiliary Text Samples
ModelGPT (Zohar et al., 2023) utilizes ChatGPT (Ouyang et al., 2022) to generate auxiliary text data by designing prompts to query ChatGPT. This extra text data mainly includes the Captions Dataset and the Synonyms Dataset.
Captions Dataset.
ModelGPT uses the following prompt to guide LLM to generate realistic and confusing text data corresponding to the user-provided classes.
Generate long and confusing image captions for the {domain} domain, which will be used to evaluate a Vision-Language Model’s {task} performance. Generate 50 captions for {classname}:
We show some generated auxiliary text examples. For example, in the category of dog, one of the text samples generated by ChatGPT is ”An adorable dog perfect for cuddles and playtime.” ModelGPT collects the results from this prompt to form the captions dataset, $\bm{D^{\text{cap}}}$.
Synonyms Dataset.
ModelGPT uses synonyms to evaluate VLM’s text encoder. For example, we expect an excellent VLM to extract similar embeddings for the words “chair” and “seat”. The prompt to guide LLM to generate synonyms is as follows.
Please list the superclasses/synonyms for {classname}. For example: chair: [furniture, seat, bench, armchair, sofa] {classname}:
ModelGPT collects the results from this prompt to form the synonyms dataset, $\bm{D^{\text{syn}}}$.
B.2 Text-Derived Scores
ModelGPT uses six metrics for model selection, which can be divided into Text Classification scores and Dataset Granularity scores. Text Classification scores include the Text top-1 accuracy score and Text f1-score. While Granularity scores include the Fisher criterion, Silhouette score, Class Dispersion score and Synonym Consistency score. Here we focus on introducing the various metrics included in the Granularity scores. We refer to the relevant content in LOVM.
Fisher Criterion $\phi_{\text{fisher}}$. The Fisher score measures the closeness of these class prompt embeddings to one another. Equation 24 shows the calculation process of it where $\hat{\bm{t}}_{i}$ is the class prompt embedding derived using the prompt ensemble strategies proposed in (Radford et al., 2021) for class $c_{i}$, $\theta(\cdot,\cdot)$ is a function that calculates the cosine similarity between two vectors, and $|C|$ is the number of classes.
$\phi_{\text{fisher}}=\frac{1}{|C|}\sum_{j=1}^{|C|}\text{max}_{i,i\neq j}\left[\theta(\hat{\bm{t}}_{i},\hat{\bm{t}}_{j})\right],$ (24)
Silhouette Score $\varphi_{\text{sil}}$. The Silhouette Score measures the separation of different-class samples in the caption dataset $\bm{D^{\text{cap}}}$. To calculate it, ModelGPT averages the cosine similarity of captions to the nearest other class by:
$\varphi_{\text{sil}}=\frac{1}{|C|}\sum_{j=1}^{|C|}\text{max}_{i,i\neq j}\left[\frac{1}{N}\sum_{k=1}^{N}\theta(\bm{D^{\text{cap}}}[j]_{k},\hat{\bm{t}}_{i})\right],$ (25)
where $\hat{\bm{t}}_{i}$ is the class prompt embedding derived using the prompt ensemble strategies proposed in (Radford et al., 2021) for class $i$, $\theta(\cdot,\cdot)$ is a function that calculates the cosine similarity between two vectors, and $|C|$ is the number of classes. $\bm{D^{\text{cap}}}[j]_{k}$ representing sample $k$ of class $j$ in the caption dataset $\bm{D^{\text{cap}}}$. There is a total of $N$ such samples for each class.
Class Dispersion Score $\rho_{\text{disp}}$. Class Dispersion Score quantifies the degree of same-class tightness or data cone radius, which is calculated using the following Equation:
$\rho_{\text{disp}}=\frac{1}{|C|N}\sum_{i=1}^{|C|}\sum_{k=1}^{N}\theta(\bm{D^{\text{cap}}}[i]_{k},\hat{\bm{t}}_{i}),$ (26)
where $\hat{\bm{t}}_{i}$ is the class prompt embedding derived using the prompt ensemble strategies proposed in Radford et al. (2021) for class $i$, $\theta(\cdot,\cdot)$ is a function that calculates the cosine similarity between two vectors, and $|C|$ is the number of classes. $\bm{D^{\text{cap}}}[i]_{k}$ representing sample $k$ of class $i$ in the caption dataset $\bm{D^{\text{cap}}}$. There is a total of $N$ such samples for each class.
Synonym Consistency Score $\gamma_{\text{syn}}$. Synonym consistency allows us to evaluate the degree of content shift between the VLMs’ pre-training and target dataset. The calculation process is shown as follows:
$\gamma_{\text{syn}}=\frac{1}{|C|N}\sum_{i=1}^{|C|}\sum_{k=1}^{N}\theta(\bm{D^{\text{syn}}}[i]_{k},\hat{\bm{t}}_{i}),$ (27)
where $\hat{\bm{t}}_{i}$ is the class prompt embedding derived using the prompt ensemble strategies proposed in (Radford et al., 2021) for class $i$, $\theta(\cdot,\cdot)$ is a function that calculates the cosine similarity between two vectors, and $|C|$ is the number of classes. $\bm{D^{\text{syn}}}[i]_{k}$ representing sample $k$ of class $i$ in the synonym dataset $\bm{D^{\text{syn}}}$. There is a total of $N$ such samples for each class.
Appendix C Implementation Details of Swab
In this section, we provide some details on the implementation of Swab, which are not mentioned in the main text due to space constraints, as well as details of some experiments related to Swab.
C.1 Filtering the Open-Source Tasks’ Classes
When the number of classes $|C_{\mathcal{S}}|$ in open-source datasets $\mathcal{S}$ is large, solving the optimal transport problem in Equation 11 can be time-consuming (as current optimal transport toolkits generally compute via CPU). To reduce the runtime of optimal transport, we can first filter the classes $C_{\mathcal{S}}$. Consider that only statistics of classes relevant to the target dataset are helpful. Therefore, we can filter out the classes in $C_{\mathcal{S}}$ that are irrelevant to the target dataset $\mathcal{T}$ based on the class-level textual semantic similarity between the open-source datasets’ classes and the target dataset’s classes. This process is shown in the following Equation:
${\bm{S}}_{ij}=\frac{\phi(c_{i}^{\mathcal{S}})^{\top}\phi(c_{j}^{\mathcal{T}})}{\|\phi(c_{i}^{\mathcal{S}})\|\cdot\|\phi(c_{j}^{\mathcal{T}})\|}\;.$ (28)
$C_{\mathcal{S}}^{{}^{\prime}}=\{c_{i}^{\mathcal{S}}|\max(\bm{S}_{i,:})>\lambda\},\ |C_{s}^{{}^{\prime}}|=k^{\prime}_{\mathcal{S}}.$ (29)
Here $\bm{S}_{i,:}$ refers to the $i$-th row of the semantic similarity matrix calculated using Equation 28, which represents the vector formed by the similarity between the $i$-th class $c_{i}^{\mathcal{S}}$ in $C_{\mathcal{S}}$ and each class $C_{\mathcal{T}}=\{c_{1}^{\mathcal{T}},\cdots,c_{k_{\mathcal{T}}}^{\mathcal{T}}\}$ of the target task. $\lambda$ is a threshold and we set $\lambda=0.5$. $k^{\prime}_{\mathcal{S}}$ refers to the number of classes in the filtered set $C_{\mathcal{S}}^{{}^{\prime}}$. Then we use the filter classes $C_{\mathcal{S}}^{{}^{\prime}}$ to calculate the transport matrix $\gamma^{*}\in\mathbb{R}^{k^{\prime}_{\mathcal{S}}\times k_{\mathcal{T}}}$ and continue with the following steps.
C.2 Using Partial Optimal Transport for Bridging the Capacity Gap
Partial optimal transport extends the optimal transport framework, enabling the selective transfer of elements from a source to a target distribution, rather than moving all elements. Its optimization problem is defined as in Equation 30. Here $z$ refers to the total amount of mass actually be transferred.
$\bm{\gamma}^{*}=\mathop{\rm argmin}_{\bm{\gamma}\in\mathbb{R}_{+}^{k^{\prime}_{\mathcal{S}}\times k_{\mathcal{T}}}}\sum_{i,j}\gamma_{i,j}\ \text{cost}_{i,j}$ (30)
[…]
$\quad\quad\bm{1}^{T}\bm{\gamma}^{T}\bm{1}=z\leq\min\left\{\|\bm{u}\|_{1},\|\bm{v}\|_{1}\right\}.$
We found that when using Equation 18 to bridge the Capacity Gap, the transport matrix $\bm{\gamma}^{*}$ obtained using partial optimal transport yields better results than the one obtained using the original optimal transport via solving the Equation 11. Therefore, in our implementation, we use the transport matrix derived from partial optimal transport to bridge the capacity gap. This also indicates that when estimating VLM’s statistics on the target dataset, different types of statistics have different preferences for the estimation methods used. This variability is worth further investigation.
Figure 5: Comparison of the consistency metrics between the accuracy calculated using text data before and after bridging the gap and the model’s true accuracy. After bridging the modality gap, the text data act as better image proxies to evaluate the model’s performance.
C.3 Data Normalization in Bridging the Modality Gap.
When bridging the Modality Gap as described in Section 3.2, we find that applying z-score normalization to the text and image features used in this process yields better results. Therefore, in our implementation, we normalize the features of all text and image samples during the modality bridging process using the following Equation:
$\bm{z}=\frac{\bm{x}-\bm{\mu}}{\bm{\sigma}}.$ (31)
Here $\bm{x}\in\mathbb{R}^{d}$ represents the image sample’s or text sample’s feature, while $\bm{\mu}\in\mathbb{R}^{d}$ and $\bm{\sigma}\in\mathbb{R}^{d}$ are calculated using the features of all samples of the same modality within its respective dataset. Besides, we also use z-score normalization to text data’s feature when we implement ModelGPT.
C.4 Task Groupings for Spider Chart Construction.
The task grouping in section 4.4 are as follows:
  • Medical: [Diabetic Retinopathy, PCam]
  • Remote: [EuroSAT, RESISC45]
  • Animals: [Oxford Pets]
  • Scene: [SUN397]
  • Traffic: [Stanford Cars, GTSRB, KITTI]
  • Universal: [ImageNet, STL10, VOC2007, CIFAR100]
C.5 Pseudo Code of Swab
Algorithm 1 shows the pseudo code of Swab.
Algorithm 1 Swab
1:  Input: Target dataset’s class names $C_{\mathcal{T}}$, open-source datasets’ class names $C_{\mathcal{S}}$, open-source datasets’ images $D^{I}_{\mathcal{S}}$.
2:  Use ChatGPT to generate auxiliary text data $\tilde{D}_{\mathcal{S}}$ and $\tilde{D}_{\mathcal{T}}$ based on $C_{\mathcal{S}}$ and $C_{\mathcal{T}}$.
3:  Calculate VLM’s class-level zero-shot image classification rankings $\{r_{m,i}\}_{m=1,i=1}^{m=M,i=k_{\mathcal{S}}}$ and class-level gap vectors $\{\bm{g}_{m,i}\}_{m=1,i=1}^{m=M,i=k_{\mathcal{S}}}$.
4:  for $k_{\mathcal{S}}$ datasets do
5:     Calculate textual similarity between the current dataset’s class names and other open-source datasets’ class names to construct a cost matrix.
6:     Solve Optimal Transport Problem based on the cost matrix to get transport matrix $\bm{\gamma}^{*}$.
7:     Use $\bm{\gamma}^{*}$ and other open-source datasets’ class-level gap vectors to predict the current dataset’s class-level gap vectors. Add the predicted class-level gap vectors to the corresponding text data’s feature of $\tilde{D}_{\mathcal{S}}$ to get modified text data.
8:  end for
9:  Calculate textual similarity between open-source datasets’ class names $C_{\mathcal{S}}$ and target dataset’s class names $C_{\mathcal{T}}$ to construct cost matrix.
10:  Solve Optimal Transport Problem based on the cost matrix to get transport matrix $\bm{\gamma}^{*}$.
11:  Use $\bm{\gamma}^{*}$ and $\{\bm{g}_{m,i}\}_{m=1,i=1}^{m=M,i=k_{\mathcal{S}}}$ to predict the class-level vectors $\{\bm{\hat{g}}_{m,i}\}_{m=1,i=1}^{m=M,i=k_{\mathcal{T}}}$ of the target dataset. Add $\{\bm{\hat{g}}_{m,i}\}_{m=1,i=1}^{m=M,i=k_{\mathcal{T}}}$ to corresponding text data’s feature of $\tilde{D}_{\mathcal{T}}$ to get modified text data.
12:  Use open-source datasets’ modified text data and images $D_{\mathcal{S}}^{I}$ to train the ranker model $f_{m}$.
13:  Use the ranker model $f_{m}$ to predict VLMs’ rankings $\{\hat{r}^{1}_{m,\mathcal{T}}\}_{m=1}^{m=M}$ based on the target datasets’ modified text data.
14:  Use $\bm{\gamma}^{*}$ and $\{\bm{g}_{m,i}\}_{m=1,i=1}^{m=M,i=k_{\mathcal{S}}}$ to predict the class-level zero-shot image classification rankings $\{\bm{\hat{r}}_{m,i}\}_{m=1,i=1}^{m=M,i=k_{\mathcal{T}}}$ of the target dataset. Calculate the average rankings $\{\hat{r}^{2}_{m,\mathcal{T}}\}_{m=1}^{m=M}$ of $\{\bm{\hat{r}}_{m,i}\}_{m=1,i=1}^{m=M,i=k_{\mathcal{T}}}$ for each VLM.
15:  Ensemble $\{\hat{r}^{1}_{m,\mathcal{T}}\}_{m=1}^{m=M}$ and $\{\hat{r}^{2}_{m,\mathcal{T}}\}_{m=1}^{m=M}$ to get final predicted rankings $\{\hat{r}^{ens}_{m,\mathcal{T}}\}_{m=1}^{m=M}$.
16:  Select the best VLM based on $\{\hat{r}^{ens}_{m,\mathcal{T}}\}_{m=1}^{m=M}$.
Appendix D More Experiment Results
In this section, we provide more experimental results of Swab.
D.1 Bridging the Modality Gap Leads to Better Image Proxies
In Section 2.3 and Figure 2, we analyze whether generated text data can act as good image proxies. Our conclusion is that due to the Modality Gap, text samples cannot directly serve as an effective substitute for images in model evaluation. To demonstrate that our method Swab can bridge this Modality Gap and thereby make text samples a better substitute for images, we conduct the following experiment.
From the figure 5, it is evident that the predicted model accuracy calculated using the modified text samples is closer to the true model accuracy compared to that calculated with the original text samples. This suggests that bridging the Modality Gap leads to better image proxies.
We use ImageNet as our dataset. First, we employ the method introduced in subsection 3.2 to predict the gap vectors for each class of the target dataset based on gap vectors calculated on open-source datasets. Then, we add the corresponding class’s gap vectors to the generated text data used in Section 2.3 and Figure 2 to bridge the modality gap. Finally, we calculate the accuracy of different models on these modified text data. We also calculate the Kendall Rank Correlation ($\tau$, higher is better) and Mean Absolute Error (MAE, lower is better) to measure the consistency between the predicted Top-1 accuracy and the true image classification accuracy.
We compare the consistency metrics of text data and modified text data. It can be observed that the consistency metrics of modified text data are better, which proves our method can reduce the gap between the generated text data and the images.
D.2 Per-Dataset Experiment Result
We present the per-dataset performance comparison between our methods Swab and ModelGPT on various datasets of LOVM benchmark in Table 9.
Table 9: LOVM Benchmark (top-1 accuracy).We compare the per-dataset experiment result between ModelGPT and Swab.
Stanford Cars CIFAR100 CLEVR-DIST. CLEVR-COUNT Country211 Retinopathy DMLab DTD EuroSAT FER2013 Flowers102 GTSRB ImageNet KITTI MNIST PCam Oxford Pets Rendered SST2 RESISC45 STL10 SUN397 SVHN VOC2007 Mean $R_{5}$ ModelGPT 0.80 0.80 0.00 0.60 0.60 0.00 0.00 1.00 0.60 0.20 0.60 0.60 0.80 0.00 0.20 0.00 0.80 0.00 1.00 0.20 0.60 0.40 0.60 0.452 Swab 0.40 0.60 0.00 0.60 0.40 0.20 0.00 0.80 0.60 0.20 0.60 0.40 0.80 0.00 0.60 0.20 0.80 0.60 0.80 0.20 0.80 0.60 0.80 0.478 $\tau$ ModelGPT 0.33 0.67 0.00 -0.33 -0.33 0.00 0.00 -0.20 1.00 0.00 1.00 1.00 1.00 0.00 0.00 0.00 0.33 0.00 -0.20 0.00 -1.00 0.00 1.00 0.186 Swab 0.00 1.00 0.00 0.33 0.00 0.00 0.00 0.00 1.00 0.00 1.00 0.00 0.33 0.00 0.33 0.00 0.33 1.00 0.00 0.00 0.33 0.33 0.33 0.275